OpenStudy (anonymous):

Evaluate 2(3x-y)^2 when x = -2 and y= -5

jimthompson5910 (jim_thompson5910):

replace x with -2 and y with -5 then evaluate

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

I got 2

jimthompson5910 (jim_thompson5910):

2 is correct

OpenStudy (anonymous):

you have to replace in the eq the values of x and y: \[2*(3*(-2)-(-5))^2\]

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

\[2*(-6+5)^2=2*(-1)^2=2*1=2\]